# Stochastic Model Predictive Control of Charging Energy Hubs with Conformal Prediction

### v1

@@ Abstract
This paper presents an online energy management system for an energy hub where electric vehicles are charged combining on-site photovoltaic generation and battery energy storage with the power grid, with the objective to decide on the battery (dis)charging to minimize the costs of operation. To this end, we devise a scenario-based stochastic model predictive control (MPC) scheme that leverages probabilistic 24-hour-ahead forecasts of charging load, solar generation and day-ahead electricity prices to achieve a cost-optimal operation of the energy hub. The probabilistic forecasts leverage conformal prediction providing calibrated distribution-free confidence intervals starting from a machine learning model that generates no uncertainty quantification. We showcase our controller by running it over a 280-day evaluation in a closed-loop simulated environment to compare the observed cost of two scenario-based MPCs with two deterministic alternatives: a version with point forecast and a version with perfect forecast. Our results indicate that, compared to the perfect forecast implementation, our proposed scenario-based MPCs are 11% more expensive, and 1% better than their deterministic point-forecast counterpart.
@@ II Charging Energy Hub Model
Figure 1: Power flow of the Charging Energy Hub composed of a grid connection ($P_{\mathrm{g}}$), a BESS ($P_{\mathrm{b}}$), a solar plant ($P_{\mathrm{pv}}$) and the total EV charging demand ($P_{\mathrm{ev}}$).
In Fig. 1, we present the power flow of the energy hub considered. It is an EV charging station supported by PV generation and a BESS. We work in discretized time with index $k$ and time resolution $\Delta T$. At each step $k$ and for each scenario $s$, the power balance of the energy hub is satisfied
@@ III-A Probabilistic Forecasting
where $F_{0}$ is the initial prediction, $\nu$ is the learning rate, $\rho_{m}$ is the line search on each boosting round $m$, $h_{m}$ are decision trees and $N_{m}$ are the total number of boosting rounds. We train the model to minimize the absolute error ($|Y-F(x)|$) on the training set, since the cost associated to our problem is linear [8]. To avoid overfitting, we use early stopping with a validation set. We use the GradientBoostingRegressor implementation from Scikit-learn [14].
@@ III-B Feature Engineering
At each $k$ and for each scenario $s$, we predict $P_{\mathrm{ev}}$ and $P_{\mathrm{pv}}$ on a 24 h window with 15 min resolution and length $N_{\mathrm{w,q}}=96$,
$\hat{\boldsymbol{P}}_{\mathrm{ev},k}^{s}=[\hat{P}_{\mathrm{ev},k}^{s},...,\hat{P}_{\mathrm{ev},k+N_{\mathrm{w,q}}-1}^{s}]^{\top},$ (14)
$\hat{\boldsymbol{P}}_{\mathrm{pv},k}^{s}=[\hat{P}_{\mathrm{pv},k}^{s},...,\hat{P}_{\mathrm{pv},k+N_{\mathrm{w,q}}-1}^{s}]^{\top},$ (15)
@@
where $d$ is the day index ($d=\lfloor\frac{k}{96}\rfloor$) and $\mathcal{T}_{d-l}=\{k\mid 96(d-l)\leq k<96(d-l+1)\}$. Similarly, we define the one-day lag of the daily standard deviation of the power as
@@
We forecast $p_{\mathrm{el}}$, at each $k$ and for each scenario $s$, on a 24 h window with 1 h resolution and length $N_{\mathrm{w,h}}=24$
$\hat{\boldsymbol{p}}_{\mathrm{el},k}^{s}=[\hat{p}_{\mathrm{el},k}^{s},...,\hat{p}_{\mathrm{el},k+N_{\mathrm{w,h}}-1}^{s}]^{\top},$ (19)
@@
where ${\boldsymbol{p}}_{\mathrm{el},k-24}=[p_{\mathrm{el},k-24},...,p_{\mathrm{el},k-1}]^{\top}$ are the observed values required to compute the 24 h differences. We design the features to predict $p_{\mathrm{el}}^{\prime}$ following [8]. We use calendar information (month of the year, day of the week, hour of the day), a lagged daily price feature (standard deviation) and two exogenous forecasts: national electricity demand ($P_{\mathrm{load}}$) and national solar and wind generation ($P_{\mathrm{gen}}$). The lagged daily price feature is equivalent to (17). For the features with exogenous forecasts of $P_{\mathrm{load}}$ and $P_{\mathrm{gen}}$, we work with normalized 24 h differences,
$\hat{\boldsymbol{P}}_{\mathrm{exog},k}^{\prime}=\frac{\hat{\boldsymbol{P}}_{\mathrm{exog},k}-\hat{\boldsymbol{P}}_{\mathrm{exog},k-24}}{P_{\mathrm{exog},d}^{\mathrm{max}}},$ (21)
where $\hat{\boldsymbol{P}}_{\mathrm{exog},k}=[\hat{P}_{\mathrm{exog},k},...,\hat{P}_{\mathrm{exog},k+23}]^{\top}$, $\hat{\boldsymbol{P}}_{\mathrm{exog},k-24}=[\hat{P}_{\mathrm{exog},k-24},...,\hat{P}_{\mathrm{exog},k-1}]^{\top}$, $P_{\mathrm{exog},d}^{\mathrm{max}}=\max(\{P_{\mathrm{exog},k}\}_{k\in\mathcal{T}_{d}})$ and, for conciseness, subscript $\mathrm{exog}$ refers to either of the two exogenous forecasts. We assume the exogenous forecasts are available at every $k$ for all hours of the predicted window. To obtain the buying price required in (11), we convert the predicted day-ahead electricity prices into as,
@@
More details on this section are available on the public repository of the article. 111{https://github.com/diegofz/ChargingEnergyHubs_MPC}
@@ III-C Generating Scenarios
For each predicted variable, we select $n_{\mathrm{s}}=1+n_{\alpha}$ scenarios from (13), where $n_{\alpha}$ is the number of percentiles used. Then we build a stochastic tree with the $n_{s}$ scenarios from each variable, assuming independence between variables. The resulting stochastic tree has $N_{s}=n_{s}^{n_{v}}$ branches, where $n_{v}$ is the number of variables considered. The probability of each branch is $\rho_{s}$. For our case, we build a stochastic tree for the predicted variables ($n_{v}=3$) and select $n_{s}=3$ scenarios from (13): $\hat{F}^{\phi}$, the 5th and the 95th percentiles.
@@ IV Model Predictive Control
where $\mathcal{S}=\{1,...,N_{\mathrm{s}}\}$, $i$ is the time step within the receding window of length $N_{\mathrm{w}}=96$ and we use notation $i|k=k+i$ to indicate that step $i$ is conditioned on information available at step $k$. We define $\mathcal{R}$ depending on the version of the MPC. For the Stochastic MPC we set =$\mathcal{R}=\{0,...,N_{\mathrm{w}}-1\}$ and for the Recourse MPC, we relax this constraint by setting =$\mathcal{R}=\{0\}$.
@@
where =$\mathcal{I}=\{0,...,N_{\mathrm{w}}-1\}$. We then solve the second-order cone program in Problem (1) to obtain the cost-optimal control and state trajectories. As the objective function, we use the average cost of operation across the scenarios
@@
where we assume equally probable scenarios ($\rho_{s}=\frac{1}{N_{\mathrm{s}}}$).
@@
$\!\min_{\boldsymbol{P}_{\mathrm{ib}}^{s}}$ $\sum_{i=0}^{N_{\mathrm{w}}-1}J_{i|k}$
@@
$\forall i\in\mathcal{I},k\in\mathcal{K},s\in\mathcal{S}.$
For comparison, we benchmark the performance of the Stochastic and Recourse MPC described, with two alternative deterministic MPCs, one using perfect forecasts (Omniscient) and another one using point-estimate forecasts from (12) (Deterministic). The deterministic MPCs are defined by setting $N_{\mathrm{s}}=1$ in Problem (1).
@@ IV-A Simulated Environment
For realistic evaluations, we use a closed-loop implementation with a simulated environment of the energy hub using observed values of $P_{\mathrm{ev}}$, $P_{\mathrm{pv}}$ and $p_{\mathrm{el}}$, and using the model defined by (1), (2), (4), (5), (10), where there is no variable uncertainty ($N_{\mathrm{s}}=1$). At each $k$, we solve Problem (1) to obtain the optimal control trajectory $\boldsymbol{P}_{\mathrm{ib},k}^{*s}=[P_{\mathrm{ib},0|k}^{*},...,P_{\mathrm{ib},N_{\mathrm{w}}-1|k}^{*s}]^{\top}$. Afterwards, we choose the first element of the control trajectory ($P_{\mathrm{ib},0|k}^{*}$) as the action, which, neither for the Stochastic nor for the Recourse MPC, depends on $s$ due to (25). Then we obtain $P_{\mathrm{b},0|k}^{*}$ from (3) and apply it to the simulated environment. The environment returns the resulting observed $E_{\mathrm{b},1|k}$ and the observed $C_{\mathrm{el},0|k}$, based on the action applied ($P_{\mathrm{b},0|k}^{*}$), the observed $P_{\mathrm{g},0|k}$ and the observed residual load ($P_{\mathrm{ev},0|k}-P_{\mathrm{pv},0|k}$). This process is repeated until $k=N_{\mathrm{ep}}-1$, when a new episode starts, resetting $k$.
@@ V Case Study Results
We use publicly available data of $P_{\mathrm{ev}}$ [16] and $P_{\mathrm{pv}}$ [17], for which both sites are located in California, USA. For $p_{\mathrm{el}}$, we employ data for the Netherlands obtained from the European Network of Transmission System Operators for Electricity (ENTSO-E) [18]. We use historical weather forecasts from the High-Resolution Rapid Refresh model by the National Oceanic and Atmospheric Administration of the US, available in the Open-Meteo.com Weather API [19]. More specifically, we work with an energy hub with a peak total EV power $\max(P_{\mathrm{ev}})=160\,\mathrm{kW}$ and a peak PV power $\max(P_{\mathrm{pv}})=70\,\mathrm{kW}$. For the BESS, we set the operational limits $\overline{E}=90\,\mathrm{kWh}$ and $\underline{E}=10\,\mathrm{kWh}$ and we initialize the battery as $E_{\mathrm{b},0}=25\,\mathrm{kWh}$. We select 2021 data to test the accuracy and coverage of the forecasts and the performance of the controllers. For a fair comparison between different seasons, we design the test set by selecting 70 days per season from 2021. Each day selected is an evaluation episode with length $N_{\mathrm{ep}}=96$. More details on data handling are available in the public repository of the article.
@@ V-A Forecasting Accuracy and Coverage
$\mathds{1}_{\mathrm{in},i}^{\alpha}=\begin{cases}1&\text{if }Y_{i}\in\mathrm{PI}_{\alpha}\\ 0&\text{otherwise}.\\ \end{cases}$ (33)
@@
Winter 0.051 0.064 0.039
Spring 0.052 0.049 0.059
Summer 0.041 0.054 0.067
Autumn 0.049 0.062 0.228
All 0.048 0.057 0.098
When forecasting $P_{\mathrm{pv}}$, there is a seasonality of the forecasting accuracy due to changing weather conditions. In Tab. I, we see that for summer and spring days, which have more stable weather conditions in this case, the predictions are more accurate. In Fig. 2, we show an example of a 24 h-ahead prediction of $P_{\mathrm{pv}}$ for two days in spring. For $p_{\mathrm{el}}$, we see a degradation of the predicting accuracy as 2021 progresses. We explain this by a regime change in the day-ahead electricity market, driven by the increasing gas prices in 2021. This highlights the additional difficulty of forecasting $p_{\mathrm{el}}$ and points toward a model improvement by introducing additional lagged features of the price or adding a feature to explain this regime change, e.g., the daily price of gas. Furthermore, in Tab. II we see that the prediction interval of $p_{\mathrm{el}}$ also degrades towards the end of 2021 and does not correspond to $\alpha=0.1$, as opposed to $\mathrm{CPI}_{\mathrm{ev},0.1}$ and $\mathrm{CPI}_{\mathrm{pv},0.1}$. In Fig. 3, we show an example of a 24 h-ahead prediction of $p_{\mathrm{el}}$ for two days in winter.
Winter 0.90 0.88 0.83
Spring 0.89 0.95 0.69
Summer 0.93 0.94 0.66
Autumn 0.90 0.87 0.22
All 0.91 0.91 0.60
We compute 24 h-ahead predictions of $P_{\mathrm{ev}}$ recursively, since we use the lagged intraday power feature defined in (18). For this reason, in Tab. III, we compare the forecasting accuracy for different prediction horizons. As expected, for shorter horizons, we see an improved prediction accuracy for $P_{\mathrm{ev}}$ and no change of performance for predicting $P_{\mathrm{pv}}$ and $p_{\mathrm{el}}$. In Fig. 4, we show an example of a 24 h-ahead prediction of $P_{\mathrm{ev}}$ for two days in spring.
Table III: Normalized Mean Absolute Error of different horizons of prediction starting at hour 0 for $P_{\mathrm{ev}}$, $P_{\mathrm{pv}}$ and $p_{\mathrm{el}}$ during a 280-day test.
Horizon $\mathrm{nMAE}_{\mathrm{ev}}$ $\mathrm{nMAE}_{\mathrm{pv}}$ $\mathrm{nMAE}_{\mathrm{el}}$
24 h 0.049 0.057 0.098
12 h 0.042 0.057 0.098
3 h 0.033 0.057 0.098

### v2

@@ Abstract
This paper presents an online energy management system for an energy hub where electric vehicles are charged combining on-site photovoltaic generation and battery energy storage with the power grid, with the objective to decide on the battery (dis)charging to minimize the costs of operation. To this end, we devise a scenario-based stochastic model predictive control (MPC) scheme that leverages probabilistic 24-hour-ahead forecasts of charging load, solar generation and day-ahead electricity prices to achieve a cost-optimal operation of the energy hub. The probabilistic forecasts leverage conformal prediction providing calibrated distribution-free confidence intervals starting from a machine learning model that generates no uncertainty quantification. We showcase our controller by running it over a 280-day evaluation in a closed-loop simulated environment to compare the observed cost of two scenario-based MPCs with two deterministic alternatives: a version with point forecast and a version with perfect forecast. Our results indicate that, compared to the perfect forecast implementation, our proposed scenario-based MPCs are 13% more expensive, and 1% better than their deterministic point-forecast counterpart.
@@ II Charging Energy Hub Model
In Fig. 1, we present the power flow of the energy hub considered.
Figure 1: Power flow of the Charging Energy Hub.
It is composed of a grid connection ($P_{\mathrm{g}}$), a BESS ($P_{\mathrm{b}}$), a solar plant ($P_{\mathrm{pv}}$) and the total EV charging demand ($P_{\mathrm{ev}}$). We work in discretized time with index $k$ and time resolution $\Delta T$. At each step $k$ and for each scenario $s$, the power balance of the energy hub is satisfied
@@ III-A Probabilistic Forecasting
where $F_{0}$ is the initial prediction, $\nu$ is the learning rate, $\rho_{m}$ is the line search on each boosting round $m$, $h_{m}$ are decision trees and $N_{m}$ are the total number of boosting rounds. We train the model to minimize the absolute error (|YF(x)||Y-F(x)|) on the training set, since the cost associated to our problem is linear [8]. To avoid overfitting, we use early stopping with a validation set. We use the GradientBoostingRegressor implementation from Scikit-learn [14].
@@ III-B Feature Engineering
At each $k$ and for each scenario $s$, we predict $P_{\mathrm{ev}}$ and $P_{\mathrm{pv}}$ on a 24 h window with 15 min resolution and length $N_{\mathrm{wq}}=96$,
$\hat{\boldsymbol{P}}_{\mathrm{ev},k}^{s}=[\hat{P}_{\mathrm{ev},k}^{s},...,\hat{P}_{\mathrm{ev},k+N_{\mathrm{wq}}-1}^{s}]^{\top}\in\mathbb{R}^{N_{\mathrm{wq}}},$ (14)
$\hat{\boldsymbol{P}}_{\mathrm{pv},k}^{s}=[\hat{P}_{\mathrm{pv},k}^{s},...,\hat{P}_{\mathrm{pv},k+N_{\mathrm{wq}}-1}^{s}]^{\top}\in\mathbb{R}^{N_{\mathrm{wq}}},$ (15)
@@
where $d$ is the day index ($d=\lfloor\frac{k}{96}\rfloor$) and $\mathcal{T}_{d-l}=\{k\mid 96\cdot(d-l)\leq k<96\cdot(d-l+1)\}$. Similarly, we define the one-day lag of the daily standard deviation of the power as
@@
We forecast $p_{\mathrm{el}}$, at each $k$ and for each scenario $s$, on a 24 h window with 1 h resolution and length $N_{\mathrm{wh}}=24$
$\hat{\boldsymbol{p}}_{\mathrm{el},k}^{s}=[\hat{p}_{\mathrm{el},k}^{s},...,\hat{p}_{\mathrm{el},k+N_{\mathrm{wh}}-1}^{s}]^{\top}\in\mathbb{R}^{N_{\mathrm{wh}}},$ (19)
@@
where ${\boldsymbol{p}}_{\mathrm{el},k-24}=[p_{\mathrm{el},k-24},...,p_{\mathrm{el},k-1}]^{\top}\in\mathbb{R}^{N_{\mathrm{wh}}}$ are the observed values required to compute the 24 h differences. We design the features to predict $p_{\mathrm{el}}^{\prime}$ following [8]. We use calendar information (month of the year, day of the week, hour of the day), a lagged daily price feature (standard deviation) and two exogenous forecasts: national electricity demand ($P_{\mathrm{load}}$) and national solar and wind generation ($P_{\mathrm{gen}}$). The lagged daily price feature is equivalent to (17). For the features with exogenous forecasts of $P_{\mathrm{load}}$ and $P_{\mathrm{gen}}$, we work with normalized 24 h differences,
$\hat{P}_{\mathrm{exog},k}^{\prime}=\frac{\hat{P}_{\mathrm{exog},k}-\hat{P}_{\mathrm{exog},k-24}}{P_{\mathrm{exog},d}^{\mathrm{max}}},$ (21)
where $P_{\mathrm{exog},d}^{\mathrm{max}}=\max(\{P_{\mathrm{exog},k}\}_{k\in\mathcal{T}_{d}})$ and, for conciseness, subscript $\mathrm{exog}$ refers to either of the two exogenous forecasts. We assume the exogenous forecasts are available at every $k$ for all hours of the predicted window. To obtain the buying price required in (11), we convert the predicted day-ahead electricity prices as,
@@
where at each step $k$ and for each scenario $s$, the resulting vectors $\hat{\boldsymbol{p}}_{\mathrm{sell},k}^{s},\hat{\boldsymbol{p}}_{\mathrm{buy},k}^{s}\in\mathbb{R}^{N_{\mathrm{wq}}}$. More details on this section are available on the public repository of the article 111{https://github.com/diegofz/ChargingEnergyHubs_MPC}.
@@ III-C Generating Scenarios
For each predicted variable, we select $n_{s}=1+n_{\alpha}$ scenarios from (13), where $n_{\alpha}$ is the number of percentiles used. Then we build a stochastic tree with the $n_{s}$ scenarios from each variable, assuming independence between variables. The resulting stochastic tree has $N_{s}=n_{s}^{n_{v}}$ branches, where $n_{v}$ is the number of variables considered. The probability of each branch is $\rho_{s}$. For our case, we build a stochastic tree for the predicted variables ($n_{v}=3$) and select $n_{s}=3$ scenarios from (13): $\hat{F}^{\phi}$, the 5th and the 95th percentiles.
@@ IV Model Predictive Control
where $\mathcal{S}=\{1,...,N_{s}\}$, $i$ is the time step within the receding window of length $N_{\mathrm{w}}=96$ and we use notation $i|k=k+i$ to indicate that step $i$ is conditioned on information available at step $k$. We define $\mathcal{R}$ depending on the version of the MPC. For the Stochastic MPC we set $\mathcal{R}=\{0,...,N_{\mathrm{w}}-1\}$ and for the Recourse MPC, we relax this constraint by setting $\mathcal{R}=\{0\}$.
@@
where $\mathcal{I}=\{0,...,N_{\mathrm{w}}-1\}$. Similarly, we convert the time index $k$ in (1), (3), (4), (5), (6) and  (11) to $i|k$ $\forall i\in\mathcal{I}$. We then solve the second-order cone program in Problem (1) to obtain the cost-optimal control and state trajectories. As the objective function, we use the average cost of operation across the scenarios
@@
where we assume equally probable scenarios ($\rho_{s}=\frac{1}{N_{s}}$).
@@
$\!\min_{\boldsymbol{P}_{\mathrm{ib},k}}$ $\sum_{i=0}^{N_{\mathrm{w}}-1}J_{i|k}$
@@
where at each $k\in\mathcal{K}$, we solve for the decision variable $\boldsymbol{P}_{\mathrm{ib},k}\in\mathbb{R}^{N_{\mathrm{w}}\times N_{s}}$.
For comparison, we benchmark the performance of the Stochastic and Recourse MPC described, with two alternative deterministic MPCs, one using perfect forecasts (Omniscient) and another one using point-estimate forecasts from (12) (Deterministic). The deterministic MPCs are defined by setting $N_{s}=1$ in Problem (1), since there is no measure of uncertainty for the predicted variables ($n_{\alpha}=0$).
@@ IV-A Simulated Environment
For realistic evaluations, we use a closed-loop implementation with a simulated environment of the energy hub using observed values of $P_{\mathrm{ev}}$, $P_{\mathrm{pv}}$ and $p_{\mathrm{el}}$, and using the model defined by (1), (2), (4), (5), (10), where there is no variable uncertainty ($N_{s}=1$). At each $k$, we solve Problem (1) to obtain the optimal control trajectory for each scenario $s$: $\boldsymbol{P}_{\mathrm{ib},k}^{\star s}=[P_{\mathrm{ib},0|k}^{\star},...,P_{\mathrm{ib},N_{\mathrm{w}}-1|k}^{\star s}]^{\top}\in\mathbb{R}^{N_{\mathrm{w}}}$. Afterwards, we choose the first element of the control trajectory ($P_{\mathrm{ib},0|k}^{\star}$) as the action, which, neither for the Stochastic nor for the Recourse MPC, depends on $s$ due to (25). Then we obtain $P_{\mathrm{b},0|k}^{\star}$ from (3) and apply it to the simulated environment. The environment returns the resulting observed $E_{\mathrm{b},1|k}$ and the observed $C_{\mathrm{el},0|k}$, based on the action applied ($P_{\mathrm{b},0|k}^{\star}$), the observed $P_{\mathrm{g},0|k}$ and the observed residual load ($P_{\mathrm{ev},0|k}-P_{\mathrm{pv},0|k}$). This process is repeated until $k=N_{\mathrm{ep}}-1$, when a new episode starts, resetting $k$.
@@ V Case Study Results
We use publicly available data of $P_{\mathrm{ev}}$ [16] and $P_{\mathrm{pv}}$ [17], for which both sites are located in California, USA. For $p_{\mathrm{el}}$ and the $\mathrm{CO_{2}}$ emissions, we employ data for the Netherlands obtained from the European Network of Transmission System Operators for Electricity (ENTSO-E) [18]. We use historical weather forecasts from the High-Resolution Rapid Refresh model by the National Oceanic and Atmospheric Administration of the US, available in the Open-Meteo.com Weather API [19]. More specifically, we work with an energy hub with a peak total EV power $\max(P_{\mathrm{ev}})=160\,\mathrm{kW}$ and a peak PV power $\max(P_{\mathrm{pv}})=70\,\mathrm{kW}$. For the BESS, we set the operational limits $\overline{E}=90\,\mathrm{kWh}$ and $\underline{E}=10\,\mathrm{kWh}$ and we initialize the battery as $E_{\mathrm{b},0}=25\,\mathrm{kWh}$. We select 2021 data to test the accuracy and coverage of the forecasts and the performance of the controllers. For a fair comparison between different seasons, we design the test set by selecting 70 days per season from 2021. Each day selected is an evaluation episode with length $N_{\mathrm{ep}}=96$. More details on data handling are available in the public repository of the article.
@@ V-A Forecasting Accuracy and Coverage
[…]
When forecasting $P_{\mathrm{pv}}$, there is a seasonality of the forecasting accuracy due to changing weather conditions. In Tab. I, we see that for summer and spring days, which have more stable weather conditions in this case, the predictions are more accurate. In Fig. 2, we show an example of a 24 h-ahead prediction of $P_{\mathrm{pv}}$ for two days in spring. For $p_{\mathrm{el}}$, we see a degradation of the predicting accuracy as 2021 progresses, with the largest $\mathrm{nMAE}$ in the Autumn. We explain this by a regime change in the day-ahead electricity market, driven by the increasing gas prices in 2021. This highlights the additional difficulty of forecasting $p_{\mathrm{el}}$ and points toward a model improvement by introducing additional lagged features of the price or adding a feature to explain this regime change, e.g., the daily price of gas. Furthermore, in Tab. II we see that the prediction interval of $p_{\mathrm{el}}$ also degrades towards the end of 2021 and does not correspond to $\alpha=0.1$, as opposed to $\mathrm{CPI}_{\mathrm{ev},0.1}$ and $\mathrm{CPI}_{\mathrm{pv},0.1}$. In Fig. 3, we show an example of a 24 h-ahead prediction of $p_{\mathrm{el}}$ for two days in winter. We compute 24 h-ahead predictions of $P_{\mathrm{ev}}$ recursively, since we use the lagged intraday power feature defined in (18). For this reason, the forecasting accuracy of $P_{\mathrm{ev}}$ improves for shorter prediction horizons. In Fig. 4, we show an example of a 24 h-ahead prediction of $P_{\mathrm{ev}}$ for two days in spring.
Winter 0.107 0.064 0.039
Spring 0.109 0.049 0.059
Summer 0.101 0.054 0.067
Autumn 0.109 0.062 0.228
All 0.106 0.057 0.098
Winter 0.89 0.88 0.83
Spring 0.90 0.95 0.69
Summer 0.91 0.94 0.66
Autumn 0.89 0.87 0.22
All 0.90 0.91 0.60